# Solving Inverse Problems in Stochastic Self-Organising Systems through Invariant Representations

### v1

@@ Abstract
Self-organising systems demonstrate how simple local rules can generate complex stochastic patterns. Many natural systems rely on such dynamics, making self-organisation central to understanding natural complexity. A fundamental challenge in modelling such systems is solving the inverse problem: finding the unknown causal parameters from macroscopic observations. This task becomes particularly difficult when observations have a strong stochastic component, yielding diverse yet equivalent patterns. Traditional inverse methods fail in this setting, as pixel-wise metrics cannot capture feature similarities between variable outcomes. In this work, we introduce a novel inverse modelling method specifically designed to handle stochasticity in the observable space, leveraging the capacity of visual embeddings to produce robust representations that capture perceptual invariances. By mapping the pattern representations onto an invariant embedding space, we can effectively recover unknown causal parameters without the need for handcrafted objective functions or heuristics. We evaluate the method on two canonical models–a reaction-diffusion system and an agent-based model of social segregation–and show that it reliably recovers parameters despite stochasticity in the outcomes. We further apply the method to real biological patterns, highlighting its potential as a tool for both theorists and experimentalists to investigate the dynamics underlying complex stochastic pattern formation.
@@ 1 Introduction
Complex systems–such as cellular automata, reaction-diffusion, or agent-based models (ABM)–self-organise into complex dynamical patterns driven by simple rules. The conventional modelling approach of complex systems consists in making an educated guess of the local rules, simulating them–be it a partial differential equation, a cellular automaton, or an agent-based model–and observing whether the emergent patterns are compatible with the phenomena we seek to explain. This modelling approach has proven highly successful in providing mechanistic understanding of a variety of phenomena, including early work on social segregation in urban environments [1], sociology [2], financial markets dynamics [3], and ecology [4, 5]. For instance, in the context of morphogenesis [6, 7], Manukyan et al. [8] identified the specific cellular automata rule and the underlying reaction-diffusion process that give rise to the skin patterns present in ocellated lizards.
@@
Figure 1: Inverse problem diagram. Inverse modelling consists in finding the mapping from observations to their causal space (also referred to as parameter space, solution space or domain space in the literature). It presents several challenges: (i) solutions may not exist for all observations; (ii) the problem of uniqueness: multiple causes may produce identical observations; and (iii) observations may be stochastic, resulting in different observed patterns for the same model parameters. Our method seeks to address the challenge of stochasticity in the observable space.
@@
Inverse modelling formulates the problem in the reverse causal direction: “Given these macroscopic observations, what are the underlying rules that produce them?[9, 10]. An inverse model takes an observation as input and returns the model, or initial configuration, that generates the observation. See Figure 1 for a visual summary of inverse modelling and some of its difficulties.
@@
A major challenge in inverse problems is the presence of stochasticity in the observations. Natural phenomena, notably those associated with complex systems, exhibit sensitivity to initial conditions. That is, the same model, under slightly different initial conditions, evolves towards divergent patterns [11, 12, 13]. Often, these patterns share high-level features but do not match when compared pixel by pixel. They are stochastic instances of the same class of patterns; see Figure 2 for an example of stochastic patterns in self-organising systems. For instance, while every human fingerprint is unique, they all share characteristic features that make them recognisable as fingerprints. Similarly, no two leopards have identical spot motifs, yet all are immediately identifiable as leopard patterns. In mathematical terms, patterns can belong to the same class in the presence of invariances–be it translational, rotational, etc.
@@
In this work, we introduce a method that addresses the challenge of solving inverse problems in systems with stochastic observable patterns. The method operates by mapping target patterns onto an invariant embedding representation, enabling the recovery of inverse parameters in stochastic self-organising systems without the need for heuristics or hand-crafted loss functions.
@@ 2 Background
Self-organising systems
Self-organisation, understood as the spontaneous emergence of global order or coordination in a system from local interactions of its parts, is observed across different physical, biological, and social systems [14, 15, 16]. Examples range from morphogenesis, where cells divide and specialise to form complex body plans in living organisms [17, 18], to animal skin patterns Murray [12] or flocking birds [19]. In social systems, this phenomenon is demonstrated in ABMs, such as Schelling’s segregation model, where simple local rules lead to emergent segregation patterns akin to those observed in cities [1].
Self-organising systems are highly sensitive to initial conditions. This sensitivity results in stochastic emergent patterns. For example, the formation of fingerprints or animal skin patterns is driven by genetic signals, local cellular feedback, and mechanical constraints that result in unique patterns on each individual. Even slight variations in initial conditions can produce vastly different results. This property of complex systems poses a significant challenge to inverse problem-solving methods: it requires defining a loss metric that captures the equivalence of patterns with inherent variability.
@@
Solving an inverse problem consists in determining the unknown parameters, or initial conditions, of a system from observations of its outcomes [9]. Contrarily, forward problems unambiguously map inputs to outputs using the system model. Formally, solving an inverse problem requires finding $(\theta,s_{0})\in\Theta\times S$ such that $F(\theta,s_{0})=y_{\text{obs}}$, where $F:\Theta\times S\to Y$ denotes the forward model mapping parameters $\theta\in\Theta$ and initial states $s_{0}\in S$ to observed data $y_{\text{obs}}\in Y$. Inverse problems are typically ill-posed, characterised by solution non-uniqueness and sensitivity [20]. These properties link inverse problems to complex systems: small changes in their parameters or initial state can result in divergent outcomes. As such, recovering the causal parameters–i.e., solving the inverse problem–shares similar challenges to ill-posed inverse problems.
Figure 2: Stochasticity in self-organising systems. Three simulation instances of two self-organising models run with the same parameters. (Top): Reaction-diffusion model; the stochasticity results from varying initial conditions $s_{0}\sim\mathcal{S}$(Bottom): Schelling’s model segregation; the stochasticity is present in the model $F$ itself in the form of asynchronous updates. Both systems are examples of stochasticity in the observable space.
Stochasticity can be present in inverse problems in two distinct ways. Stochasticity in the causal space: multiple parameters or initial conditions $(\theta,s_{0})\in\Theta\times S$ map to identical or nearly identical observations $y_{\text{obs}}\in Y$, resulting in degenerate solutions and non-uniqueness, i.e., non-injectivity of $F$. In this case, the unknown quantities in the causal space are distributions. An example is seismic tomography, which uses seismic waves recorded at the surface to reconstruct the Earth’s interior. Solutions are non-unique: similar surface measurements can originate from different inner structures. Stochasticity in the observable space: in this case, observations are inherently stochastic. Either resulting from randomness in initial states, such as in reaction-diffusion systems ($y_{\text{obs}}=F(\theta,s_{0})$ with $s_{0}\sim\mathcal{S}$), or from intrinsic stochasticity in the forward model $F$, as in the asynchronous updates of an agent-based model. See Figure 2. The existing literature on inverse stochastic modelling exclusively focusses on the first case: how to address stochasticity in the causal space [21, 22, 23, 24, 25, 26], where similar observations can have different causal origins. On the contrary, the method proposed in this work addresses the challenge of stochasticity in the observable space.
@@
Visual embeddings are learnt vector representations that capture relevant features of images beyond the raw pixel values. Instead of encoding only low-level information, such as colour and intensity, embeddings represent higher-level visual attributes, such as shape, texture, structure, and semantic content [27, 28, 29, 30, 31]. Visual embedding models transform image data into a lower-dimensional space where semantically or visually similar images are mapped to nearby points, while dissimilar images are placed farther apart [32, 33, 34]. See diagram in Figure 3a.
Figure 3: Invariant representations. (a): Diagram representing models parameters $\theta_{i}$, generating observable patterns $y_{i}$, and the visual patterns being mapped to an embedding $z_{i}$. (b): Distance matrix between embedding representations $z$ of four patterns (two reaction-diffusion and two ABM segregation patterns). Notice that each of the patterns is an unique instantiation of the models. Similarity clusters appear, suggesting that the embedding model represent patterns with similar features as similar vectors. The values displayed are pairwise cosine distances $1-\cos(z_{i},z_{j})$.
Geometric invariant representations. Embedding representations have been shown to be invariant under different pattern transformations [35, 36, 37, 38]. For instance, Moskvyak and Maire [32] showed that convolutional embeddings can produce invariant representations of animal patterning under homographic transformations. Our method uses CLIP [33], an open-source vision and text embedding model trained on large-scale datasets using contrastive learning (in this work, only the visual encoder is used). Its embeddings, represented as $512$-dimensional vectors, capture high-level visual features. The large variety of images seen by CLIP during training results in rich representations, making them suitable for general-purpose applications without requiring additional fine-tuning. Figure 3b shows that CLIP embeddings are able to capture relevant visual features.
@@ 3 Proposed Method
A common approach to solving inverse problems is to formulate them as optimisation problems by defining a loss function that measures the discrepancy between target data and predictions [21, 22]. However, pixel-based metrics are unable to meaningfully capture feature similarities between stochastic patterns. Therefore, they cannot account for the intrinsic variations often present in self-organising patterns. To solve inverse problems with stochasticity in the observable space, a metric capable of capturing the feature-level similarities rather than exact pixel-level matches is needed. Embedding representations offer an effective solution to this challenge by encoding features and invariances of the patterns. The idea behind the proposed method is straightforward: map stochastic patterns onto an embedding space where perceptually similar patterns have similar vector representations. Then, use these invariant representations to solve the inverse problem of finding the unknown parameters that generate the stochastic patterns we seek to reconstruct. A visual summary is provided in Figure 4.
@@
Figure 4: Method overview. A target pattern is mapped onto an embedding space $z_{\text{target}}\in\mathbb{R}^{\mathcal{D}}$ using the embedding model $E$, where $D$ is the embedding space dimension. A forward model, parametrised by $\theta\in\mathbb{R}^{\mathcal{M}}$, where $\mathcal{M}$ is the degrees of freedom of the model, generates a pattern which is mapped to the same embedding space resulting in $z_{\text{predicted}}$. The loss $\mathcal{L}$ between target and predicted embedding guides the optimiser, iteratively updating $\theta$ until convergence. By operating in embedding space, the method offers robustness against the inherent stochasticity of emergent patterns.
@@
We demonstrate the method using CLIP, a pre-trained embedding model, to encode the visual patterns into vector embeddings $z$. Both the target and the predicted patterns are mapped onto the same embedding space, resulting in representations despite stochastic variations. A black-box optimiser is then used to search for the inverse parameters that minimise the embedding distance between the target pattern and the predictions.
@@
We use CMA-ES (Covariance Matrix Adaptation Evolution Strategy), a population-based optimiser [39], to search the causal space. Evolutionary Strategies algorithms have been shown to find high-performing solutions in a variety of tasks [40, 41, 42]–including in the context of inverse problems [43, 44]. CMA-ES operates by iteratively updating the covariance matrix of a multivariate normal distribution at each generation, so that candidate solutions $\theta_{i}$ are increasingly guided towards promising regions of the search space. Solutions are iteratively sampled from a Gaussian distribution centred on the current mean with covariance defining the search distribution: $\boldsymbol{\theta}_{i}=\mathbf{m}+\Delta\mathbf{C}^{1/2}\mathbf{\sigma}$ with $\mathbf{\sigma}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ where $\mathbf{m}$ is the mean, $\Delta$ is the step-size, $\mathbf{C}$ is the covariance matrix and $\mathbf{I}$ is the identity matrix. The mean $\mathbf{m}$ is updated as a weighted sum of selected solutions $\mathbf{m}_{\text{new}}=\sum_{k=1}^{\mu}w_{k}\boldsymbol{\theta}_{k}$ where $\mu$ is the number of selected solutions, and $w_{k}$ are the weights based on solution rank. The step-size $\sigma$ is adjusted to balance exploration and convergence.
@@
The method operates as follows: (1) sample a set of parameters $\theta_{i}$ and random initial states $s_{0}\sim\mathcal{S}$, (2) generate the corresponding patterns $y_{i}$ through the forward model $F$, (3) use the embedding model to encode the resulting patterns into the embedding space $z$, (4) compute the loss between the target $z_{target}$ and the candidate solutions, and (5) update the optimiser accordingly. In the case of CLIP, the loss metric is cosine similarity. By focusing on high-level visual similarities rather than pixel-level matches, the method is robust to stochastic variations, and is able to solve inverse problems across entire families of patterns with shared visual features–without requiring handcrafted loss functions. The method is detailed in Algorithm 1.
Algorithm 1 Inverse modelling through embedding representation
@@
1:Forward model $F$, State space $S$, Loss function $\mathcal{L}$, Embedding model $E$, Optimiser
@@
4:     Sample candidate solutions $\theta_{i}$ from Optimiser
@@
6:     Run simulations: $y_{\text{predicted}}\leftarrow F(\theta_{i},s_{0})$
@@
9:         $z_{\text{predicted}}\leftarrow E(y_{\text{predicted}})$
10:     Compute loss: $loss_{i}\leftarrow\mathcal{L}(z_{\text{predicted}},z_{\text{target}})$
11:     Update optimiser state with $loss_{i}$
@@
The codebase and implementation of the algorithm can be found here: https://github.com/enajx/InverseStochastic.

### v2

# Solving Inverse Problems in Stochastic Self-organizing Systems through Invariant Representations

@@ Abstract
Self-organizing systems demonstrate how simple local rules can generate complex stochastic patterns. Many natural systems rely on such dynamics, making self-organization central to understanding natural complexity. A fundamental challenge in modeling such systems is solving the inverse problem: finding the unknown causal parameters from macroscopic observations. This task becomes particularly difficult when observations have a strong stochastic component, yielding diverse yet equivalent patterns. Traditional inverse methods fail in this setting, as pixel-wise metrics cannot capture feature similarities between variable outcomes. In this work, we introduce a novel inverse modeling method specifically designed to handle stochasticity in the observable space, leveraging the capacity of visual embeddings to produce robust representations that capture perceptual invariances. By mapping the pattern representations onto an invariant embedding space, we can effectively recover unknown causal parameters without the need for handcrafted objective functions or heuristics. We evaluate the method on three self-organizing systems: a physical, a biological, and a social one; namely, a reaction-diffusion system, a model of embryonic development, and an agent-based model of social segregation. We show that the method reliably recovers parameters despite stochasticity in the pattern outcomes. We further apply the method to real biological patterns, highlighting its potential as a tool for both theorists and experimentalists to investigate the dynamics underlying complex stochastic pattern formation.
@@ 1 Introduction
Complex systems–such as cellular automata, reaction-diffusion, or agent-based models (ABM)–self-organise into complex dynamical patterns driven by simple rules. The conventional modeling approach of complex systems consists of making an educated guess of the local rules, simulating them–be it a partial differential equation, a cellular automaton, or an agent-based model–and observing whether the emergent patterns are compatible with the phenomena we seek to explain. This modeling approach has proven highly successful in providing mechanistic understanding of a variety of phenomena, including early work on social segregation in urban environments [43], sociology [3], financial markets dynamics [42], and ecology [30, 12]. For instance, in the context of morphogenesis [6, 15], Manukyan et al. [29] identified the specific cellular automata rule and the underlying reaction-diffusion process that give rise to the skin patterns present in ocellated lizards.
@@
Figure 1: Inverse problem diagram. Inverse modeling consists in finding the mapping from observations to their causal space (also referred to as parameter space, solution space or domain space in the literature). It presents several challenges: (i) solutions may not exist for all observations; (ii) the problem of uniqueness: multiple causes may produce identical observations; and (iii) observations may be stochastic, resulting in different observed patterns for the same model parameters. Our method seeks to address the challenge of stochasticity in the observable space.
@@
Inverse modeling formulates the problem in the reverse causal direction: “Given these macroscopic observations, what are the underlying rules that produce them?[44, 7]. An inverse model takes an observation as input and returns the model, or initial configuration, that generates the observation. See Fig. 1 for a visual summary of inverse modeling and some of its difficulties.
@@
A major challenge in inverse problems is the presence of stochasticity in the observations. Natural phenomena, notably those associated with complex systems, exhibit sensitivity to initial conditions. That is, the same model, under slightly different initial conditions, evolves towards divergent patterns [11, 33, 1]. Often, these patterns share high-level features but do not match when compared pixel by pixel. They are stochastic instances of the same class of patterns–see Fig. 2 for an example of stochastic patterns in self-organizing systems. For instance, while every human fingerprint is unique, they all share characteristic features that make them recognisable as fingerprints. Similarly, no two leopards have identical spot motifs, yet all are immediately identifiable as leopard patterns. In mathematical terms, patterns can belong to the same class in the presence of invariances–be it translational, rotational, or of another kind.
@@
In this work, we introduce a method that addresses the challenge of solving inverse problems in systems with stochastic observable patterns. The method operates by mapping target patterns onto an invariant embedding representation, enabling efficient recovery of inverse parameters in stochastic self-organizing systems without the need for heuristics or hand-crafted loss functions.
@@ 2 Background
Self-organizing systems
Self-organization, understood as the spontaneous emergence of global order or coordination in a system from local interactions of its parts, is observed across different physical, biological, and social systems [19, 23, 37]. Examples range from morphogenesis, where cells divide and specialise to form complex body plans in living organisms [51, 52], to animal skin patterns Murray [33] or flocking birds [40]. In social systems, this phenomenon is demonstrated in ABMs, such as Schelling’s segregation model, where simple local rules lead to emergent segregation patterns akin to those observed in cities [43].
Self-organizing systems are highly sensitive to initial conditions. This sensitivity results in stochastic emergent patterns. For example, the formation of fingerprints or animal skin patterns is driven by genetic signals, local cellular feedback, and mechanical constraints that result in unique patterns on each individual. Even slight variations in initial conditions can produce vastly different results. This property of complex systems poses a significant challenge to inverse problem-solving methods: it requires defining a loss metric that captures the equivalence of patterns with inherent variability.
@@
Solving an inverse problem consists in determining the unknown parameters, or initial conditions, of a system from observations of its outcomes [44]. By contrast, forward problems unambiguously map inputs to outputs using the system model. Formally, solving an inverse problem requires finding $(\theta,s_{0})\in\Theta\times S$ such that $F(\theta,s_{0})=y_{\text{obs}}$, where $F:\Theta\times S\to Y$ denotes the forward model mapping parameters $\theta\in\Theta$ and initial states $s_{0}\in S$ to observed data $y_{\text{obs}}\in Y$. Inverse problems are typically ill-posed, characterised by solution sensitivity and non-uniqueness [18]. These properties link inverse problems to complex systems: small changes in their parameters or initial state can result in divergent outcomes. As such, recovering the causal parameters–i.e., solving the inverse problem–shares similar challenges with ill-posed problems.
Figure 2: Stochasticity in self-organizing systems. Three simulation instances of two self-organizing models run with the same parameters. (Top): Reaction-diffusion model; the stochasticity results from varying initial conditions $s_{0}\sim\mathcal{S}$(Bottom): Schelling’s model segregation; the stochasticity is present in the model $F$ itself in the form of asynchronous updates. Both systems are examples of stochasticity in the observable space.
Stochasticity can be present in inverse problems in two distinct ways. Stochasticity in the causal space: multiple parameters or initial conditions $(\theta,s_{0})\in\Theta\times S$ map to identical or nearly identical observations $y_{\text{obs}}\in Y$, resulting in degenerate solutions and non-uniqueness, i.e., non-injectivity of $F$. In this case, the unknown quantities in the causal space are distributions. An example is seismic tomography, which uses seismic waves recorded at the surface to reconstruct the Earth’s interior. Solutions are non-unique: similar surface measurements can originate from different inner structures. Stochasticity in the observable space: in this case, observations are inherently stochastic. Either resulting from randomness in initial states, such as in reaction-diffusion systems ($y_{\text{obs}}=F(\theta,s_{0})$ with $s_{0}\sim\mathcal{S}$), or from intrinsic stochasticity in the forward model $F$, as in the asynchronous updates of an agent-based model. See Fig. 2. The existing literature on inverse stochastic modeling exclusively focuses on the first case: how to address stochasticity in the causal space [49, 54, 27, 24, 53, 14], where similar observations can have different causal origins. By contrast, the method proposed in this work addresses the challenge of stochasticity in the observable space.
@@
Visual embeddings are learned vector representations that capture relevant features of images beyond the raw pixel values. Instead of encoding only low-level information, such as colour and intensity, embeddings represent higher-level visual attributes, such as shape, texture, structure, and semantic content [13, 35, 22, 34, 45]. Visual embedding models transform image data into a lower-dimensional space where semantically or visually similar images are mapped to nearby points, while dissimilar images are placed farther apart [32, 39, 25]. See diagram in Fig. 3a.
Figure 3: Invariant representations. (a): Diagram representing models parameters $\theta_{i}$, generating observable patterns $y_{i}$, and the visual patterns being mapped to an embedding $z_{i}$. (b): Distance matrix between embedding representations $z$ of four patterns (two reaction-diffusion and two ABM segregation patterns). Notice that each of the patterns is a unique instantiation of the model. Similarity clusters appear, suggesting that the embedding model represents patterns with similar features as similar vectors. The values displayed are pairwise cosine distances $1-\cos(z_{i},z_{j})$.
Geometric invariant representations. Embedding representations have been shown to be invariant under different pattern transformations [16, 56, 46, 55, 21]. For instance, Moskvyak and Maire [32] showed that convolutional embeddings can produce invariant representations of animal patterning under homographic transformations. We demonstrate our method with CLIP [39], an open-source vision and text embedding model trained on large-scale datasets using contrastive learning (in this work, only the visual encoder is used). Its embeddings, represented as $512$-dimensional vectors, capture high-level visual features. The large variety of images seen by CLIP during training results in rich representations, making them suitable for general-purpose applications without requiring additional fine-tuning. Fig. 3b shows that CLIP embeddings are able to capture relevant visual features.
@@ 3 Proposed Method
A common approach to solving inverse problems is to formulate them as optimization problems by defining a loss function that measures the discrepancy between target data and predictions [49, 54]. However, pixel-based metrics are unable to meaningfully capture feature similarities between stochastic patterns. Therefore, they cannot account for the intrinsic variations often present in self-organizing patterns. To solve inverse problems with stochasticity in the observable space, a metric capable of capturing the feature-level similarities rather than exact pixel-level matches is needed. Embedding representations offer an effective solution to this challenge by encoding features and invariances of the patterns. The idea behind the proposed method is straightforward: map stochastic patterns onto an embedding space where perceptually similar patterns have similar vector representations. Then, use these invariant representations to solve the inverse problem of finding the unknown parameters that generate the stochastic patterns we seek to reconstruct. A visual summary is provided in Fig. 4.
@@
Figure 4: Method overview. A target pattern is mapped onto an embedding space $z_{\text{target}}\in\mathbb{R}^{\mathcal{D}}$ using the embedding model $E$, where $D$ is the embedding space dimension. A forward model, parametrised by $\theta\in\mathbb{R}^{\mathcal{M}}$, where $\mathcal{M}$ is the degrees of freedom of the model, generates a pattern which is mapped to the same embedding space resulting in $z_{\text{solution}}$. The loss $\mathcal{L}$ between the target and the candidate solution embedding guides the optimizer, iteratively updating $\theta$ until convergence. By operating in embedding space, the method offers robustness against the inherent stochasticity of emergent patterns.
@@
We demonstrate the method using CLIP, a pre-trained embedding model, to encode the visual patterns into vector embeddings $z$. Both the target and the predicted patterns are mapped onto the same embedding space, resulting in similar representations despite stochastic variations. A black-box optimizer is then used to search for the inverse parameters that minimise the embedding distance between the target pattern and the predictions.
@@
We use CMA-ES (Covariance Matrix Adaptation Evolution Strategy), a population-based optimizer [20], to search the causal space. Evolutionary strategies have been shown to find high-performing solutions in a variety of tasks [41, 2, 48]–including in the context of inverse problems [5, 17]. CMA-ES operates by iteratively updating the covariance matrix of a multivariate normal distribution at each generation, so that candidate solutions $\theta_{i}$ are increasingly guided towards promising regions of the search space. Solutions are iteratively sampled from a Gaussian distribution centred on the current mean with the covariance defining the search distribution: $\boldsymbol{\theta}_{i}=\mathbf{m}+\Delta\mathbf{C}^{1/2}\mathbf{\sigma}$ with $\mathbf{\sigma}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ where $\mathbf{m}$ is the mean, $\Delta$ is the step-size, $\mathbf{C}$ is the covariance matrix and $\mathbf{I}$ is the identity matrix. The mean $\mathbf{m}$ is updated as a weighted sum of selected solutions $\mathbf{m}_{\text{new}}=\sum_{k=1}^{\mu}w_{k}\boldsymbol{\theta}_{k}$ where $\mu$ is the number of selected solutions, and $w_{k}$ are the weights based on solution rank.
@@
The method operates as follows: (1) sample a set of parameters $\theta_{i}$ and random initial states $s_{0}\sim\mathcal{S}$, (2) generate the corresponding patterns $y_{i}$ through the forward model $F$, (3) use the embedding model to encode the resulting patterns into the embedding space $z$, (4) compute the loss between the target $z_{target}$ and the candidate solutions’ embeddings $z_{solution}$, and (5) update the optimizer accordingly. In the case of CLIP, the loss metric is cosine similarity. By focusing on high-level visual similarities rather than pixel-level matches, the method is robust to stochastic variations, and is able to solve inverse problems across entire families of patterns with shared visual features–without requiring handcrafted loss functions. The method is detailed in Algorithm 1.
Algorithm 1 Inverse modeling through embedding representation
@@
1:Forward model $F$, State space $S$, Loss function $\mathcal{L}$, Embedding model $E$, optimizer
@@
4:  Sample candidate solutions $\theta_{i}$ from optimizer
@@
6:  Run simulations: $y_{\text{solution}}\leftarrow F(\theta_{i},s_{0})$
@@
9:    $z_{\text{solution}}\leftarrow E(y_{\text{solution}})$
10:  Compute loss: $loss_{i}\leftarrow\mathcal{L}(z_{\text{solution}},z_{\text{target}})$
11:  Update optimizer state with $loss_{i}$